Xyp+y^2q+2x-xyz

1
Expert's answer
2021-11-17T06:42:51-0500

Solution;

"xyp+y^2q+2x^2-xyz"

Rewrite as;

"xyp+y^2q=xyz-2x^2"

Here;

"P=xy"

"Q=y^2"

"R=xyz-2x^2"

Axillary equations are;

"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"

"\frac{dx}{xy}=\frac{dy}{y^2}=\frac{dz}{xyz-2x^2}" ......(1)

Taking the first two fractions in (1)

"\frac{dx}{xy}=\frac{dy}{y^2}"

Simplify;

"\frac{dx}{x}=\frac{dy}{y}"

Integrating respectively;

"log(x)=log(y)+log(c_1)"

Therefore;

"log(\frac xy)=log(c_1)\implies\frac xy=_-^+c_1=k" .....(2)

Taking the last two equations in (1);

"\frac{dy}{y^2}=\frac{dz}{xyz-2x^2}"

Substituting (2);

"\frac{dy}{y^2}=\frac{dz}{ky^2z-2k^2y^2}"

Simplify as;

"kdy=\frac{1}{z-2k}dz"

Integrating ;

"ky=log(z-2k)+c_2"

Using (2) as substitution;

"x-log(z-\frac{2x}{y})=c_2" ....(3)

The general solution is given by (2) and (3);

"(\frac xy,x-log(z-\frac{2x}{y})=0"